Can someone please help me with this?


Given P(A) = .40, P(B) = .50. If A and B are independent, find P(A∩B).

Answer by stanbon(75887)

Given P(A) = .40, P(B) = .50. If A and B are independent, find P(A∩B).
---------------------------------
P(A and B) = P(A|B)*P(B)
---
If you assume that A and B are independent you get:
P(A and B) = P(A)*P(B) = 0.4*0.5 = 0.2
------------
If you cannot assume that A and B are independent
you don't have enough information to determine
P(A and B).
================
Cheers,
Stan H.